OpenStudy (anonymous):

Write the slope-intercept inequality for the graph below. If necessary, use <= for or >= for

OpenStudy (calculusxy):

Where is the graph?

OpenStudy (anonymous):

y-y1<m(x-x1) ?

OpenStudy (calculusxy):

the equation is y=mx+b.

OpenStudy (calculusxy):

so just substitute the letters for the numbers

OpenStudy (anonymous):

0=4(-1)+b?

OpenStudy (anonymous):

0=-4+b?

OpenStudy (anonymous):

b=4

OpenStudy (anonymous):

Okay well first off do you know how to find slope?

OpenStudy (anonymous):

slope 4

OpenStudy (anonymous):

correct. do you know the point-slope form?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay so this is where its probably going to be important to know that formula. its: y-y1=m(x-x1)

OpenStudy (anonymous):

so basically now you plug and chug

OpenStudy (anonymous):

so pick a point, im going to pick (0,2) and then put it into the equation. you then get: y-2=4(x-0). now you solve for y.

OpenStudy (anonymous):

y+2=4(x+1)

OpenStudy (anonymous):

that works!

OpenStudy (anonymous):

solve for y now :)

OpenStudy (anonymous):

y=4x+2

OpenStudy (anonymous):

correct! so now you have your equation, all you need to do is just figure out if you need a less than or equal to, or, a greater than or equal to symbol

OpenStudy (anonymous):

y<4x+2?

OpenStudy (anonymous):

<=

OpenStudy (anonymous):

y<=4x+2?

OpenStudy (anonymous):

give me one second to clarify

OpenStudy (anonymous):

okay yes i agree! :) do you understand how i got the equation?